OpenStudy (anonymous):

which equation has a semetry of x=1/5? y = −5x2 − x + 2 y = −5x2 + x + 2 y = −5x2 + 2x − 9 y = −5x2 − 2x − 9

OpenStudy (amoodarya):

turn them to standard form

OpenStudy (anonymous):

vertex must be 1/5. ok??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

y=ax^2+bx+c vertex=-b/a follow that.

OpenStudy (anonymous):

uh

OpenStudy (amoodarya):

y = −5x2 + 2x − 9 this ok but you try to compute !

OpenStudy (anonymous):

? i am confused

OpenStudy (anonymous):

help

OpenStudy (amoodarya):

look at image you find out